OpenStudy (anonymous):

There are 270 girls and 210 boys enrolled in the high school. The cover of the year book will show a photograph of a pair of students. How many pairs are possible?

OpenStudy (jamanorthi):

add 270 and 210 whatever u get divide that number by 2 and u will get the answer

OpenStudy (debbieg):

Not quite. There are 480 total students. So the question is, how many unique pairs can be selected from those 480 students. That's a combination, 480C2 (since the same 2 students are the same pair, whatever order they are selected). Do you know the formula for that?

OpenStudy (debbieg):

\[nCr=\frac{ n! }{ r!(n-r)! }\]

OpenStudy (anonymous):

Jamanorthi's answer is wrong. That high school has 270+210=480 students. You can chose 480 different students from all students. After that you can chose 479 students, because you have already selected one of them. That means 480*479 ordered pairs are possible. But order is not important so you should divide that number by 2! . The answer is 480*479/2= 114960. (If the pair in cover book consists of one male and one female student the answer is 270*210=56700.)